# Classification of Histopathology Slides with Persistence Homology Convolutions

### v1

@@ 5 Results
For each of the five data representations — persistent homology convolutions, global persistent homology, thresholded images, and combinations thereof — we chose the model hyperparemeters that resulted in the highest accuracy. Performance metrics for these accuracy-maximizing models are summarized in Table 1. The models trained with local persistent homology the form of PHCs exhibit better performance on all metrics than the models trained with either grayscale images or global persistence. This supports our hypotheses that the local arrangement of topological features is an important characteristic for the differentiation of the three tissue classes, and that at least some of this information is not detected by models trained with a standard CNN architecture.
We also tested the performance of PHC computing using two other filtrations: the alpha complex on pixels of thresholded images and the lower-star filtration on the magnitude of pixels in $X$ as used in [30]. The accuracy maximizing hyperparameters were chosen using the same pipeline as above, however, Step 3 from 4.1.2 is omitted when computing lower-star filtration and persistence is instead computed directly on the thresholded image. Their performance is summarized in Table 2. It is apparent from Table 2 that models trained with the lower-star filtration are inadequate at distinguishing between histopathological classes, performing no better than random. Models trained with height PHC data tend to yield the best performance across all metrics with the exception of sensitivity and specificity where it either ties with the alpha complex or slightly under-performs. These results indicate that the information encoded by the height PHC — namely the location and linear size of cells — are especially important for histopathological classification in this dataset.
@@
Table 3 compares the runtimes required to compute and process PHC and global persistent homology for this dataset across three different filtrations. This includes image preprocessing (thresholding), computing the persistent homology data (either local or global), and vectorizing the resulting data. As expected, persistent homology convolutions can be computed much more quickly than global persistent homology.
@@
We also consider how model performance varies across the hyperparameter sweep. Tables 4 and 5 display the average metric performance and standard error of each data type aggregated across all trained models. Models trained with PHC data generated using height PH or the alpha complex have higher average scores and less variance for each performance metric. Interestingly, the height PHC data has higher average scores in accuracy and precision whereas the alpha complex PHC data has higher average scores in sensitivity and specificity. This indicates that models trained with PHCs on specific filtrations may exhibit less sensitivity to the choice of parameters and may therefore be easier to work with in practice. Moreover, it suggests that PHC can extract meaningful geometric information from the micrographs.
@@
Global Height PH 0.7424 0.7600 0.6943 0.5437
Height PHC 0.9170 0.9167 0.9389 0.9585
Image Data 0.9083 0.9119 0.9432 0.9563
Images + Global PH 0.9171 0.9170 0.9563 0.9607
Images + Height PHC 0.9389 0.9430 0.9651 0.9716
Table 1: Evaluation metrics for the accuracy-maximizing models for each data type. The best performing data types for each metric are highlighted in bold.
Filtration Type Accuracy Precision Sensitivity Specificity
Alpha PHC 0.9170 0.9170 0.9432 0.9585
Height PHC 0.9170 0.9167 0.9389 0.9585
Lowerstar PHC 0.4280 0.4928 0.2576 0.1485
Images + Alpha PHC 0.9345 0.9427 0.9651 0.9804
Images + Height PHC 0.9389 0.9430 0.9651 0.9716
Images + Lowerstar PHC 0.4672 0.5225 0.2489 0.1834
Filtration Type Alpha Height Lower Star
PHC Data 2.55 sec 11.35 sec 1.09 sec
Global PH Data 6.80 sec 3496.0 sec 2.70 sec
Table 3: Average run time to compute persistence on a single image across various complex types on a sample of a $N=100$ images from the Ost. dataset. Computations were performed on an M1 MacBook Pro. The fastest compute time is highlighted in bold.
Data Representation Accuracy Precision Sensitivity Specificity
Global Height PH $0.6300\pm 0.0709$ $0.6758\pm 0.0758$ $0.4694\pm 0.1019$ $0.4631\pm 0.0825$
Height PHC $0.8480\pm 0.0412$ $0.8528\pm 0.0385$ $0.8720\pm 0.0638$ $0.8609\pm 0.0828$
Image Data $0.8051\pm 0.0935$ $0.8101\pm 0.0941$ $0.7987\pm 0.1674$ $0.7063\pm 0.3015$
Images + Global Height PH $0.8250\pm 0.0808$ $0.8291\pm 0.0807$ $0.8287\pm 0.1510$ $0.7657\pm 0.2703$
Images + Height PHC $\boldsymbol{0.8653\pm 0.0561}$ $\boldsymbol{0.8695\pm 0.0523}$ $\boldsymbol{0.8905\pm 0.0964}$ $\boldsymbol{0.8748\pm 0.1667}$
Table 4: Average evaluation metric across all 1000 model initializations. The best performing data types for each metric are highlighted in bold. Observe that the Height PHC + Images have the highest averages and Height PHC has the smallest standard error.
Filtration Type Accuracy Precision Sensitivity Specificity
Alpha PHC $0.8597\pm 0.0221$ $0.8624\pm 0.0220$ $\boldsymbol{0.8914\pm 0.0301}$ $\boldsymbol{0.8853\pm 0.0517}$
Height PHC $0.8480\pm 0.0412$ $0.8528\pm 0.0385$ $0.8720\pm 0.0638$ $0.8609\pm 0.0828$
Lower Star PHC $0.3338\pm 0.0378$ $0.3425\pm 0.1098$ $0.1098\pm 0.0787$ $0.0783\pm$ 0.0869
Images + Alpha PHC $0.8316\pm 0.0848$ $0.8379\pm 0.0811$ $0.8288\pm 0.1774$ $0.7637\pm 0.2934$
Images + Height PHC $\boldsymbol{0.8653\pm 0.0561}$ $\boldsymbol{0.8695\pm 0.0523}$ $0.8905\pm 0.0964$ $0.8748\pm 0.1667$
Images + Lower Star PHC $0.3342\pm 0.0273$ $0.3342\pm 0.0273$ $0.1076\pm 0.0790$ $0.0021\pm 0.0089$
Table 5: Average evaluation metric across all 1000 model initializations across multiple filtrations. The best performing data types for each metric are highlighted in bold. Observe that the PHC + Images have the highest averages and PHC has the smallest standard error using height PH and the alpha complex.
@@ Acknowledgments
The authors are grateful for the support of the National Science Foundation under Award No. 2232967. Additionally, we would like to thank Ajay Krishna Vajjala, Deval Parikh, Tyrus Berry, Anthony Pizzimenti, and Jeremy Mason for their thoughtful comments and suggestions on the content of this paper.

### v3

@@ 5 Results
For each of the five data representations — persistent homology convolutions, global persistent homology, truncated images, and combinations thereof — we chose the model hyperparemeters that resulted in the highest accuracy. A variety of filtration methods were tested to measure the performance of PHC. The full set of results are available in the Appendix Acknowledgments. Here, we restrict our focus to the two best performing filtrations in dimension one: the alpha complex and extended persistence on the adjacency complex. Performance metrics for these accuracy-maximizing models are summarized in Table 1. The models trained with local persistent homology the form of PHCs using the alpha complex and extended persistence of the adjacency complex exhibit better performance on all metrics than the models trained with either grayscale images or global persistence. This supports our hypotheses that the local arrangement of topological features is an important characteristic for the differentiation of the three tissue classes, and that at least some of this information is not detected by models trained with a standard CNN architecture.
Note that in some instances, specific evaluation scores for the alpha complex and the extended adjacency complex are identical. There are two factors that likely contribute to this outcome. First, the repetition in scores is an upper limit of model performance over the hyperparameter sweep; the repetition in scores were present among several of the top performing models across the sweeps. Second, the image dataset from (Arunachalam et al., 2019) are subdivided images of whole slides. In some instances, a subdivision of whole slides yields images with an absence of tissue matter, potentially obfuscating training. This may be reflected in the evaluation score with models repeatedly misclassifying samples.
Models trained with adjacency XPHC data tend to yield the best performance across all metrics with the exception of sensitivity and specificity where it either ties with the alpha complex or slightly under-performs. That said, those two filtrations yield very similar performance despite encoding qualitatively different information about the geometry of histopathology slides. Namely, the alpha complex encodes the geometry of the cellular and subcelluar structure whereas the adjacency complex measures the grayscale value of cells in a locality. Moreover, the improved performance of models trained with PHC indicates that additional information encoded — namely the locality and cell geometry — are important for histopathological classification in this dataset.
@@
Table 2 compares the runtimes required to compute and process PHC and global persistent homology for this dataset across the filtrations. This includes image preprocessing (thresholding/truncating), computing the persistent homology data (either local or global), and vectorizing the resulting data. As expected, persistent homology convolutions can be computed much more quickly than global persistent homology.
@@
We also consider how model performance varies across the hyperparameter sweep. Table 4 display the average metric performance and standard error of each data type aggregated across all trained models. Models trained with PHC data using any complex have higher average scores and less variance for each performance metric. Interestingly, the adjacency XPHC data has higher average scores in accuracy and precision whereas the alpha complex PHC data has higher average scores in sensitivity and specificity. This indicates that models trained with PHCs on specific filtrations may exhibit less sensitivity to the choice of parameters and may therefore be easier to work with in practice. Moreover, it suggests that PHC can extract meaningful geometric information from the micrographs.
Further evidence for this hypothesis is found by examining the intrinsic dimension estimates of the various data representations presented in Table 3. The intrinsic dimension of the PHC of the data is substantially lower than that of the full images. This indicates that PHC produces a simplified geometric summary that retains information important to the underlying biology. On the other hand, global PH loses even more information and leads to worse performance. This suggests the importance of making the correct choice of geometric summary, and in particular that the locality of geometric features is significant for histopathological classification in this data set.
Filtration Type Accuracy Precision Sensitivity Specificity
Grayscale Image Data 90.8% 91.2% 94.3% 95.6%
Alpha Complex PHC 91.7% 91.7% 94.3% 95.9%
Adjacency Complex XPHC 91.7% 91.7% 93.9& 95.9%
Global Adjacency Complex XPH 74.2% 76.0% 69.4% 54.4%
Images + Global Adjacency Complex XPH 91.7% 91.7% 95.6% 96.1%
Images + Alpha Complex PHC 93.5% 94.3% 96.5% 98.0%
Images + Adjacency Complex XPHC 93.9% 94.3% 96.5% 97.2%
Filtration Type PHC Data Global PH Data
Alpha Complex 2.6 sec 6.8 sec
Extended Adjacency Complex 11.4 sec 3496.0 sec
Table 2: Average run time to compute persistence on a single image across various complex types on a sample of a $N=100$ images from the Ost. dataset. The dimension of homology computed is fixed to $p=1$. Computations were performed on an M1 MacBook Pro. The fastest compute time is highlighted in bold.
Data Representation Global Local
Alpha Complex PH 0.18 4.41
Adjacency Complex XPH 0.18 2.24
Histopathology Images 24.67 12.50
Table 3: Intrinsic dimension estimates of various data representations computed using the maximum likelihood estimator proposed by (Levina and Bickel, 2004; MacKay, David and Ghahramani, Zoubin, 2005). These results were computed with a sample size of $N=500$.
@@
Global Adjacency Complex XPH $63.0\pm 7.1\%$ $67.6\pm 7.6\%$ $46.9\pm 10.19\%$ $46.3\pm 8.3\%$
Adjacency Complex XPHC $84.8\pm 4.1\%$ $85.3\pm 3.9\%$ $87.2\pm 6.4\%$ $86.1\pm 8.3\%$
Alpha Complex PHC $86.0\pm 2.2\%$ $86.2\pm 2.2$ $\boldsymbol{89.1\pm 3.0}$ $\boldsymbol{88.5\pm 5.2}$
Image Data $80.5\pm 9.4\%$ $81.0\pm 9.4\%$ $79.9\pm 16.7\%$ $70.6\pm 30.2\%$
Images + Alpha Complex PHC $83.2\pm 8.5\%$ $83.8\pm 8.1\%$ $82.9\pm 17.7\%$ $76.4\pm 29.3\%$
Images + Global Adjacency Complex XPH $82.5\pm 8.1\%$ $82.9\pm 8.1\%$ $82.9\pm 15.1\%$ $76.6\pm 27.0\%$
Images + Adjacency Complex XPHC $\boldsymbol{86.5\pm 5.6\%}$ $\boldsymbol{87.0\pm 5.2\%}$ $89.1\pm 9.6\%$ $87.5\pm 16.7\%$
Table 4: Average evaluation metric across all 1000 model initializations. The best performing data types for each metric are highlighted in bold. Observe that the Lowerstar XPHC + Images have the highest averages in accuracy and precision and the TPHC has the highest average sensitivity and specificity scores.
@@ Acknowledgments
The authors are grateful for the support of the National Science Foundation under Award No. 2232967. Additionally, we would like to thank Ajay Krishna Vajjala, Deval Parikh, and Anthony Pizzimenti for their suggestions regarding the implementation of PHC. The authors would also like to thank Tyrus Berry, Heather Harrington, Kelly Maggs, and Jeremy Mason for their thoughtful comments regarding the content of this paper. Computational resources were provided by the Office of Research Computing at George Mason University.
Appendix A Full PHC Experiments
To evaluate the performance of PHC, we test five filtrations: the ordinary and extended persistence of the alpha complex and the lower star filtration, and the extended persistence of the adjacency complex [^3] Ordinary persistence was not computed of the adjacency complex as it contains information equivalent to that of the extended persistence. The lower star filtration was computed on a simplicial complex obtained by subdividing the cubical complex formed by the pixels. Our experimental setup and computational framework is publicly available in our repository: https://github.com/Shrunalp/PHC.git. We compute the persistent homology of each filtration in dimensions zero and one. For each filtration, we investigate six training configurations determined by whether the filtration is inputted together with image data or alone, and whether 0-dimensional features, $1$-dimensional features, or both are included. To accommodate these diverse inputs, the CNN architecture utilizes separate convolution and pooling branches for each data format.
Tables 5-9 summarize the accuracy-maximizing models across all filtration. Hybrid models, trained using image and PHC data, consistently outperform the benchmarks of models trained exclusively on image data (shown in Table 1). Furthermore, while extended persistence yields better results than ordinary persistence for the lower star filtration, it underperforms for the alpha complex.
Furthermore, we observe that the best performance was given by the one-dimensional persistent homology features, except for the extended alpha complex. This suggests that the geometry of the cells themselves (e.g. the irregularity of cell shape and size, the presence of multinucleation, etc) is more relevant to tumor classification than the geometry of the empty space between cells.
Observe that models trained on the persistent homology of the adjacency complex perform better than those trained on the persistent homology of the lower star filtration. This is counterintuitive, as the information contained by the former is also encoded in the latter by construction. We hypothesize that the geometric features represented by the adjacency complex are more relevant to the classification than the additional features encoded by the lower star filtration, and that including more information than necessary — even if it meaningfully corresponds to geometric features — results in worse model performance. This could also explain the underperformance of extended persistence of the alpha complex as compared to the ordinary persistence of the same filtration.
Lastly, Table 10 compares the run times of PHC relative to global PH computations across filtrations. As expected, PHC computations are significantly faster.
Alpha Complex Dimension Model Performance
Test Accuracy Precision Sensitivity Specificity
With Images 0 90.0% 90.4% 93.9% 94.3%
1 93.5% 94.3% 96.5% 98.0%
0 + 1 91.7% 92.8% 95.2% 96.5%
Without Images 0 83.8% 85.5% 86.9% 83.2%
1 91.7% 91.7% 94.4% 95.9%
0 + 1 87.2% 87.6% 89.5% 87.8%
Table 5: Evaluation metrics for the accuracy-maximizing models trained using Alpha Complex PHC data. Models are trained across different dimensions in homology, with and without image data.
Extended Persistence Dimension Model Performance
Alpha Complex Test Accuracy Precision Sensitivity Specificity
With Images 0 88.7% 90.0% 90.8% 91.3%
1 90.0% 89.9% 93.0% 93.2%
0 + 1 91.7% 91.7% 95.2% 96.1%
Without Images 0 81.2% 83.3% 83.0% 80.6%
1 83.8% 87.3% 87.7% 88.9%
0 + 1 86.0% 89.8% 87.3% 84.9%
Table 6: Evaluation metrics for the accuracy-maximizing models trained using Extended Persistence on the Alpha Complex PHC data. Models are trained across different dimensions in homology, with and without image data.
Lower Star Dimension Model Performance
Filtration Test Accuracy Precision Sensitivity Specificity
With Images 0 92.1% 92.1% 94.3% 96.5%
1 92.1% 93.0% 94.3% 96.5%
0 + 1 90.8% 90.8% 95.2% 95.4%
Without Images 0 91.3% 92.0% 95.2% 96.1%
1 88.7% 88.6% 90.4% 90.8%
0 + 1 90.4% 90.3% 91.7% 95.2%
Table 7: Evaluation metrics for the accuracy-maximizing models trained using Lower Star Filtration PHC data. Models are trained across different dimensions in homology, with and without image data.
Extended Persistence Dimension Model Performance
Lower Star Test Accuracy Precision Sensitivity Specificity
With Images 0 92.1% 91.7% 94.3% 96.5%
1 92.6% 92.5% 94.3% 96.3%
0 + 1 90.8% 91.2% 93.8% 96.1%
Without Images 0 91.2% 92.0% 95.2% 96.1%
1 90.0% 90.3% 95.2% 95.0%
0 + 1 90.3% 90.3% 91.7% 95.1%
Table 8: Evaluation metrics for the accuracy-maximizing models trained using Extended Lower Star Filtration PHC data. Models are trained across different dimensions in homology, with and without image data.
Extended Persistence Dimension Model Performance
Adjacency Complex Test Accuracy Precision Sensitivity Specificity
With Images 0 88.2% 88.9% 92.1% 92.6%
1 93.9% 94.3% 96.5% 97.2%
0 + 1 91.7% 92.8% 95.2% 96.5%
Without Images 0 83.0% 82.6% 83.4% 77.1%
1 91.7% 91.7% 93.9% 95.9%
0 + 1 90.0% 90.0% 91.7% 94.1%
Table 9: Evaluation metrics for the accuracy-maximizing models trained using Extended Persistence of the Adjacency Complex PHC data. Models are trained across different dimensions in homology, with and without image data.
Filtration Type PHC Data Global PH Data
Alpha Complex 2.6 sec 6.8 sec
Lower Star Filtration 5.6 sec 73.4 sec
Extended Adjacency Complex 11.4 sec 3496.0 sec
Extended Alpha Complex 8.1 sec 10.6 sec
Extended Lower Star Filtration 16.2 sec 92.9 sec
Table 10: Average run time to compute persistence on a single image across various complex types on a sample of a $N=100$ images from the Ost. dataset. The dimension of homology computed is fixed to $p=1$. Computations were performed on an M1 MacBook Pro. The fastest compute time is highlighted in bold.